OpenStudy (anonymous):

et f(x) = x^2-1 and g(x) = 2x-3 Find (g o f)(-2)

OpenStudy (anonymous):

Sorry, I meant what's g(x^2-1) And then plug in -2.

OpenStudy (anonymous):

= 2(x^2)-1 g(f(-2)) = 2((-2)^2)-1 8-1 = 7